OpenStudy (tia_sweetheart17):

solve the equation. 9-3x-6-3x=5

OpenStudy (anonymous):

\[9-3x-6-3x=5\]\[9-6x-6=5\]\[3-6x=5\]\[-6x=2\]\[x=-1/3\] \[9-3(-1/3)-6-3(-1/3)=5\]\[9+1-6+1=5\]\[10-6+1=5\]